# Non-Markovian Quantum Jump Method for Driven-Dissipative Two-Level Systems

### v1

@@ Abstract
We propose a modified non-Markovian quantum jump method to overcome the obstacle of exponentially increased trajectory number in conventional quantum trajectory simulations. In our method the trajectories are classified into the trajectory classes characterized by the number of quantum jumps. We derive the expression of the existence probability of each trajectory (class) which is essential to construct the density matrix of the open quantum system. This modified method costs less computational resource and is more efficient than the conventional quantum trajectory approach. As applications we investigate the dynamics of spin-1/2 systems subject to Lorentzian reservoirs with considering only the no-jump and one-jump trajectories. The revival of coherence and entanglement induced by the memory effect is observed.
@@ I Introduction
Quantum system cannot be completely isolated, it always inevitably interacts with its surrounding environment. The interaction between the system and the environment leads the evolution of the state of quantum system to be non-unitary which is different from the case of a closed quantum system breuer_book . To describe the dynamics of open quantum systems, the Markovian approximation is usually employed, which assumes that the evolution of the system’s state depends only on its current state without reference to its history lindblad1976 ; gks1976 . In the Markovian dynamics the information of the systems flows unidirectionally from the system to the environment without any feedback breuer_PRL2009 ; breuer_RMP2016 . A powerful and efficient approach for simulating the Markovian process is the quantum jump method which considers the stochastic time-evolution of a large number of pure states plenio_RMP1998 ; daley2014 . Comparing with the simulation of the full density matrix via the quantum master equation in the Lindblad form, such unraveled process requires less memory since it only manipulates the state vector in each realization donvil_NC2022 . This has made the Markovian quantum jump method a standard tool in investigating the dynamics of open quantum systems.
@@
The non-Markovian time-evolution of the state of an open quantum system can be described by the master equation with memory kernel shabani2005 ; ciccarello2013 . The memory kernel is an integral function about the memory time and weights the past state in the master equation. Besides, the time-local Lindblad master equation can also describe the memory effects in the time-evolution. When the system of interest interacts with a structured reservoir, i.e. the spectral density of varies appreciably with the frequency of the environmental modes, the decay rate becomes time-dependent and even to be negative temporarily haikka_PRA2010 or permanently megier_SR2017 . The negative decay rate may give rise to the possibility for the system to recover the state that before decoherence. In the framework of Markovian quantum jump method, this can be implemented by performing the reversed quantum jumps to cancel the normal jumps previously occurred when the decay rate was positive. This idea has been put forward in the non-Markovian quantum jump (NMQJ) approach which basically unravels the non-Markovian time-local master equation piiloPRL2008 .
@@
Despite the significant progress has been made in applying NMQJ to single-qubit systempiiloPRL2008 ; piiloPRA2009 ; becker_PRL2023 , it is still rather demanding to extend this method to the quantum many-body systems chiriaco_PRB2023 . The complexity of quantum many-body systems arises not only from internal quantum entanglement and interactions, but also from more intricate interactions between the system and its environment. In general, the system’s state after a normal quantum jump is not the eigenstate of the effective non-Hermitian Hamiltonian of the open system and may evolve to excited state that is feasible for next quantum jump. Therefore the number of quantum trajectories in the many-body NMQJ simulation grows exponentially with the size of the system. Furthermore, the calculation of the transition probabilities between different trajectories become quite involved and the conservation of the total probability may be broken due to the presence of the reversal quantum jump. Alternatively, introducing an external driving on a single quantum system also presents the challenges similar to those in many-body system, as the number of quantum trajectories likewise grows exponentially. This exponential growth not only complicates the probability of reverse jumps in the negative decay rate region but also significantly increases computational complexity, making the NMQJ method difficult to apply in such cases.
@@
This paper is organized as follows. In Sec. II we introduce the modified NMQJ method by starting with a brief review of the method of quantum jump in Markovian and Non-Markovian cases. The concepts of quantum trajectory and trajectory class are introduced in Sec. II.3 which are essential in our method. The existence probabilities of each quantum trajectory in the Markovian and non-Markovian cases are derived in Sec. II.4 thus enabling the calculation of the expectation value of the observable. In Sec. III, we demonstrate two toy models of spin-1/2 systems that we are going to apply the NMQJ method. In Sec. IV we investigate the dynamics of the systems. In particular we concentrate on how the non-Markovianity affects the recover of coherence in the single-spin system as well as the nonmonotonic entanglement in two-spin system. We summarize in Sec. V.
@@ II.3 Quantum trajectory and trajectory class
In order to implement the NMQJ method in a realistic numerical simulation, the time axis is discretized by short interval $\delta t$ as shown in the Fig. 1. The time interval is so short that only one quantum jump could take place within $\delta t$. Suppose that the initial state of the system is $|\psi(t_{0})\rangle$. The blocks in the same color constitute a specific evolution path of the system’ state, which we refer to as a quantum trajectory, denoted by $H_{n}^{\alpha}$. In addition, in a certain quantum trajectory, the quantum state of the system at time $t_{s}$ is denoted by $|\psi_{n}^{\alpha}(t_{s})\rangle$. The superscript $\alpha$ is an array of time and represents the time sequence that the normal quantum jump occurred. The subscript $n=\#\alpha$ represents the number of the normal jumps. A quantum trajectory describes the continuous evolution of the system’s state solely governed by the effective non-Hermitian Hamiltonian $\hat{H}_{\text{eff}}$ from the latest time point in the time sequence, i.e. $t=\alpha(n)$. Specially, in Fig. 1 the quantum trajectory in blue describes the evolution of the system’s state without any quantum jumps, denoted by $H_{n}^{\o}$ where $\o$ denotes the empty time sequence. The state of $\hat{H}_{n}^{\o}$ at time $t_{s}$ is thus given by
$|\psi_{0}^{\o}(t_{s})\rangle=\frac{\exp{(-i\hat{H}_{\text{eff}}t_{s})}|\psi(t_{0})\rangle}{||\exp{(-i\hat{H}_{\text{eff}}t_{s})}|\psi(t_{0})\rangle||},$ (12)
@@ II.4 Existence probability of quantum trajectory
Suppose that the total number of realizations is $N$ and the number of realizations stay in the trajectory $H_{n}^{\alpha}$ at time $t_{w}$ is $N_{n}^{\alpha}(t_{w})$. In the limit of infinite $N$, the existence probability of each trajectory is defined as
@@ II.4.1 $\Delta(t)>0$ case
Let us first focus on the time duration $[t_{0},t_{P}]$ in which the decay rate is positive $\Delta(t)>0$. To start, we employ the result reported in Ref. piiloPRL2008 , the existence probability of quantum trajectory (class) $H_{0}^{\o}$ at $t_{w}\in[t_{0},t_{P}]$ is given by
@@
which is the probability of a quantum jump occurring at $t_{w}$ conditioned on that no jump takes place before $t_{w}$ piiloPRA2009 . In Eq. (17) the time sequence has been expressed by the single-element array $[t_{w}]$ for clarity.
@@
with $\alpha=[\alpha^{\prime},t_{w}]$. $K_{n}^{\alpha}(t_{P})$ is the product of the existence probability of a quantum trajectory underwent $(n-1)$ quantum jumps before $t_{w-1}$, the probability of a quantum jump takes place at $t_{w}$, and the probability of no more jump takes place in the future $t>t_{w}$. The probability of a quantum jump occurs at $t$ with respect to the state $|\psi_{n}^{\alpha}(t)\rangle$ is
@@ II.4.2 $\Delta(t)<0$ case
stands for the probability that a two-jump trajectory $H_{2}^{\alpha^{\prime}}$ may jump back to $H_{1}^{\alpha}$ at $t_{P}$, which depends on the population with respect to the state of the target trajectory $|\psi_{1}^{\alpha}(t_{P})\rangle$. Eq. (31) possesses a quite intuitive physical meaning: The first term of the r.h.s. of Eq. (31) represents the reduction of $K_{1}^{\alpha}(t_{P})$ due to the reversed jump to its mother trajectory $H_{0}^{\o}$, while the second term represents the increment of $K_{1}^{\alpha}(t_{P})$ due to the contributions via the reversed jump of all the trajectories belonging to class $\{H_{2}^{\alpha^{\prime}}\}$ Substituting Eqs. (23) and (32) into Eq. (31), one can obtain
@@
Moreover the calculation can be truncated at a certain hierarchy $n^{*}$ if the existence probability of trajectory class $\{H_{n}^{\alpha}\}$ with $n>n^{*}$ is below a threshold. Indeed, as will be seen soon, the quantum trajectory $H_{0}^{\o}$ and the trajectory class $\{H_{1}^{\alpha}\}$ occupy almost all the portion of the possible quantum trajectories in a realistic example. Therefore, we may safely neglect the contributions of the trajectories those undergo more than two quantum jumps. The computational cost in storing the history of time-evolution is dramatically reduced.
@@ III The models
[…]
@@
[…]
@@
$J_{\text{Lor}}(\nu)=\frac{\alpha^{2}}{2\pi}\frac{\Gamma^{2}}{(\nu-\omega)^{2}+\Gamma^{2}}.$ (43)
When characteristic time scale $\tau_{S}$ is much shorter than the reservoir correlation time $\tau_{C}$, one can make the secular approximation on the system-environment interaction, leading to the following time-dependent decay rate in the time-local Lindblad master equations (41) and (42),
@@
where $q_{0}$ denotes the detuning of the central frequency of the Lorentzian spectrum to the frequency of two-level system, and $\eta$ denotes the coupling strength between the system and the environment. Here it should be pointed out that in the rest of the paper we work in units of $\Gamma^{-1}$ and set $\eta=10$ and $q_{0}=6$ in the numerical computation. The time interval is chosen as $\delta t=10^{-3}\Gamma^{-1}$.
@@ IV.1 Model I: revival of coherence
[…]
@@
[…]
@@
[…]
@@ IV.2 Model II: Sudden death and revival of entanglement
V Summary

### v2

@@ Abstract
We propose a modified non-Markovian quantum jump method to overcome the obstacle of dramatically increased trajectory number in conventional quantum trajectory simulations. In our method the trajectories are classified into the trajectory classes characterized by the number of quantum jumps. We derive the expression of the existence probability of each trajectory (class), which is essential to construct the density matrix of the open quantum system. This modified method costs less computational resources and is more efficient than the conventional quantum trajectory approach. As applications we investigate the dynamics of spin-1/2 systems subject to Lorentzian reservoirs with considering only the no-jump and one-jump trajectories. The revival of coherence and entanglement induced by the memory effect is observed.
@@ I Introduction
Quantum system cannot be completely isolated, it always inevitably interacts with its surrounding environment. The interaction between the system and the environment leads the evolution of the state of quantum system to be non-unitary which is different from the case of a closed quantum system breuer_book . To describe the dynamics of open quantum systems, the Markovian approximation is usually employed, which assumes that the evolution of the system’s state depends only on its current state without reference to its history lindblad1976 ; gks1976 . In the Markovian dynamics the information of the systems flows unidirectionally from the system to the environment without any feedback breuer_PRL2009 ; breuer_RMP2016 . A powerful and efficient approach for simulating the Markovian process is the quantum jump method which considers the stochastic time-evolution of a large number of pure states plenio_RMP1998 ; daley2014 . Comparing with the simulation of the full density matrix via the quantum master equation in the Lindblad form, such unraveled process requires less memory since it only manipulates the state vector in each realization Dalibard1992 ; Molmer1993 . This has made the Markovian quantum jump method a standard tool in investigating the dynamics of open quantum systems.
@@
The non-Markovian time-evolution of the state of an open quantum system can be described by the master equation with memory kernel shabani2005 ; ciccarello2013 . The memory kernel is an integral function about the memory time and weights the past state in the master equation. Besides, the time-local Lindblad master equation can also describe the memory effects in the time-evolution. When the system of interest interacts with a structured reservoir, i.e. the spectral density of varies appreciably with the frequency of the environmental modes, the decay rate becomes time-dependent and even to be negative temporarily haikka_PRA2010 or permanently megier_SR2017 . In this work, we will consider the quantum spin interacts with a Lorentzian reservoir. In the limit of weak coupling between the spin and the environment, the time-convolutionless projection technique yields the time-local master equation in Lindblad form describing the time-evolution of the spin under study. In particular, the non-Markovian feature of such master equation is indicated by the dissipator with negative decay rate haikka_PRA2010 ; haikka_PhysScr_2010 . The negative decay rate may give rise to the possibility for the system to recover the state that before decoherence. In the framework of Markovian quantum jump method, this can be implemented by performing the reversed quantum jumps to cancel the normal jumps previously occurred when the decay rate was positive. This idea has been put forward in the non-Markovian quantum jump (NMQJ) approach which basically unravels the non-Markovian time-local master equation piiloPRL2008 .
@@
Despite the significant progress has been made in applying NMQJ to single-qubit system piiloPRL2008 ; piiloPRA2009 ; smirne_PRL2020 ; chruscinski_quantum2022 ; donvil_NC2022 ; becker_PRL2023 ; settimo_PRA2024 , it is still rather demanding to extend this method to the quantum many-body systems chiriaco_PRB2023 . The complexity of quantum many-body systems arises not only from internal quantum entanglement and interactions, but also from more intricate interactions between the system and its environment. In general, the system’s state after a normal quantum jump is not the eigenstate of the effective non-Hermitian Hamiltonian of the open system and may evolve to excited state that is feasible for next quantum jump. Therefore the number of the possible path of the time-evolution in many-body NMQJ simulation grows dramatically. Because after a normal quantum jump acting locally on the $n$-th subsystem, the state of the system is a product state of the ground state of the $n$-th subsystem and a state of the rest of the joint system. The normalization coefficient of such state depends on the non-zero amplitude of the state at the moment when the normal quantum jump occurred and thus gives rise to different initial states for the following time-evolution. As a consequence, a large number of time-evolution paths need to be recorded compared with the single-body case. Furthermore, the calculation of the transition probabilities between different trajectories become quite involved and the conservation of the total probability may be broken due to the presence of the reversal quantum jump. Alternatively, introducing an external driving on a single quantum system also presents the challenges similar to those in many-body system, as the number of quantum trajectories likewise grows dramatically. This significant growth not only complicates the probability of reverse jumps in the negative decay rate region but also significantly increases computational complexity, making the NMQJ method difficult to apply in such cases.
@@
This paper is organized as follows. In Sec. II we introduce the modified NMQJ method by starting with a brief review of the method of quantum jump in Markovian and Non-Markovian cases. The concepts of quantum trajectory and trajectory class are introduced in Sec. II.3 which are essential in our method. The existence probabilities of each quantum trajectory in the Markovian and non-Markovian cases are derived in Sec. II.4 thus enabling the calculation of the expectation value of the observable. In Sec. III, we demonstrate two toy models of spin-1/2 systems that we are going to apply the NMQJ method. In Sec. IV we investigate the dynamics of the systems. In particular we concentrate on how the non-Markovianity affects the recover of coherence in the single-spin system as well as the nonmonotonic entanglement in two-spin system. We summarize in Sec. VI.
@@ II.3 Quantum trajectory and trajectory class
[…]
@@ II.4 Existence probability of quantum trajectory
Suppose that the total number of realizations is $N$ and the number of realizations that stay in the trajectory $H_{n}^{\alpha}$ at time $t_{w}$ is $N_{n}^{\alpha}(t_{w})$. In the limit of infinite $N$, the existence probability of each trajectory is defined as
@@ II.4.1 $\Delta(t)>0$ case
Let us first focus on the time duration $t_{0},t_{P}]$ in which the decay rate is positive $\Delta(t)>0$. To start, we employ the result reported in Ref. piiloPRL2008 , the existence probability of quantum trajectory (class) $H_{0}^{\o }$ at $t_{w}\in[t_{0},t_{P}]$ is given by
@@
which is the probability of a quantum jump occurring at $t_{w}$ conditioned on that no jump takes place before $t_{w}$ piiloPRA2009 . In Eq. (17) the time sequence has been expressed by the single-element array $t_{w}]$ for clarity.
@@
with $\alpha=[\alpha^{\prime},t_{w}]$. $K_{n}^{\alpha}(t_{P})$ is the product of the existence probability of a quantum trajectory underwent (n1)(n-1) quantum jumps before $t_{w-1}$, the probability of a quantum jump takes place at $t_{w}$, and the probability of no more jump takes place in the future $t>t_{w}$. The probability of a quantum jump occurs at $t$ with respect to the state $|\psi_{n}^{\alpha}(t)\rangle$ is
@@ II.4.2 $\Delta(t)<0$ case
stands for the probability that a two-jump trajectory $H_{2}^{\alpha^{\prime}}$ may jump back to $H_{1}^{\alpha}$ at $t_{P}$, which depends on the population with respect to the state of the target trajectory $|\psi_{1}^{\alpha}(t_{P})\rangle$. Eq. (31) possesses a quite intuitive physical meaning: The first term of the r.h.s. of Eq. (31) represents the reduction of $K_{1}^{\alpha}(t_{P})$ due to the reversed jump to its mother trajectory $H_{0}^{\o }$, while the second term represents the increment of $K_{1}^{\alpha}(t_{P})$ due to the contributions via the reversed jump of all the trajectories belonging to class $\{H_{2}^{\alpha^{\prime}}\}$ . Substituting Eqs. (23) and (32) into Eq. (31), one can obtain
@@
Moreover the calculation can be truncated at a certain hierarchy $n^{*}$ if the existence probability of trajectory class $\{H_{n}^{\alpha}\}$ with $n>n^{*}$ is below a threshold. Indeed, as will be seen soon, in the presented models in Sec. III, the quantum trajectory $H_{0}^{\o }$ and the trajectory class $\{H_{1}^{\alpha}\}$ occupy almost all the portion of the possible quantum trajectories in a realistic example. Therefore, in such situations, we may safely neglect the contributions of the trajectories those undergo more than two quantum jumps. The computational cost in storing the history of time-evolution is dramatically reduced. If the existence probability of the quantum trajectory that experiences multiple quantum jumps is non-negligible, the contribution must be taken into account in the computation and therefore a higher truncation number $n^{*}$ is needed.
@@ III The models
In order to check the performance, we are going to apply the modified NMQJ method to two toy models. We restrict the discussion on the dynamics of spin-1/2 systems (with two levels $|\uparrow\rangle$ and $|\downarrow\rangle$) subject to noisy environment as shown in Fig. 2. Here we assume that the temperature of the environment is zero such that the thermal excitation can be neglected luoma_PRL2020 . Thus the environment can only induce the deexcitation of the spin through the incoherent flip down to the $z$-direction.
@@
where $\hat{H}_{\text{I}}=\Omega\hat{\sigma}^{x}$ presents an external driving field imposing on the spin along $x$-direction, $\Omega$ is the Rabi frequency, $\hat{\sigma}^{\alpha}$ with $\alpha=x,y,z$ are Pauli matrices for spin-1/2 system and the raising and lowering operators are $\hat{\sigma}^{\pm}\equiv(\hat{\sigma}^{x}\pm i\hat{\sigma}^{y})/2$. The notation $\{\cdot,\cdot\}$ stands for the anti-commutator.
@@
$J_{\text{Lor}}(\nu)=\frac{\eta^{2}}{2\pi}\frac{\Gamma^{2}}{(\nu-\omega)^{2}+\Gamma^{2}},$ (43)
where $\eta$ denotes the coupling strength between the system and the environment. The spectral density of the quantized electromagnetic field inside an imperfect cavity is well approximated by the the spectrum in Eq. (43). The width of the distribution quantifying the leakage of photons through the cavity mirrors is characterized by the parameter $\Gamma$. When characteristic time scale $\tau_{S}$ is much shorter than the reservoir correlation time $\tau_{C}$, one can make the secular approximation on the system-environment interaction, leading to the following time-dependent decay rate in the time-local Lindblad master equations (41) and (42),
@@
where $q_{0}$ denotes the detuning of the central frequency of the Lorentzian spectrum to the frequency of two-level system haikka_PhysScr_2010 . The limit of weak coupling is reached when $\eta^{2}$ is smaller than the smallest relevant frequency in the system. Here in the rest of the paper we work in units of $\Gamma$ and set $\eta=10$ and $q_{0}=6$ in the numerical computation. The time intervals in both models are chosen as $\delta t=10^{-3}\Gamma^{-1}$.
@@ IV.1 Model I: revival of coherence
Let us first investigate the dynamics of the single spin-1/2 system in Model I. The Rabi frequency of the driving field is chosen as $\Omega=0.5$. The spin precesses around the external field along $x$-direction which leads to a coherent oscillation between the $|\uparrow\rangle$ and $|\downarrow\rangle$ states. The normal quantum jump always incoherently flips the spin to the state $|\downarrow\rangle$. The jump probability is proportional to the population of the $|\uparrow\rangle$ state. Therefore the existence probability of the quantum trajectories depends on the initial state of the spin. We characterize the initial state via the polar and azimuthal angles $(\theta,\phi)$ in the Bloch sphere.
@@
In order to determine the truncated number of quantum jumps, we check the the total existence of probability of quantum trajectories undergoes two quantum jumps at the end of the positive decay region. The value of $\sum_{\alpha}{K_{2}^{\alpha}(t_{P})}$ with various initial states are shown in Fig. 4. For the initial states on the equator, the maximal $K_{2}(t_{P})$ locates at $\phi=\frac{\pi}{2}$ corresponding to the spin polarizing along the positive $y$-direction. This is because the precession brings the Bloch vector passing by the north pole leading to more probability of normal quantum jump. Oppositely the initial state along negative $y$-direction will result in less $K_{2}(t_{P})$ because the precession is counterclockwise, both the coherent and incoherent spin-flipping tend to reduce the population of $|\uparrow\rangle$ state thus suppress the probability of normal quantum jump. It is also worth noting that the magnitude of $K_{2}(t_{P})$ in the whole $\theta-\phi$ plane is less than $10^{-2}$, indicating the contribution of the two-jump trajectories is so small that can be neglected in the computation.
In Fig. 5 more details of the time-evolution of the single spin for $t\leq t_{N}$ are shown with various initial states: the spin-up state, spin-down state and the superposition of both. As mentioned above, since the $K_{2}$ is very small we only consider the trajectories $H_{0}^{\o }$ and $\{H_{1}^{\alpha}\}$. The upper panels of Fig. 5 shows the time-dependence of $K_{0}^{\o }(t)$ and the total existence probability of all the trajectories with one-jump $K_{1}(t)=\sum_{\alpha}{K_{1}^{\alpha}(t)}$. In the Markovian region (blue shaded), because the normal jumps occur the $K_{0}^{\o }$ decreases monotonically while $K_{1}(t)$ increases. Moreover, from Figs. 5(a)-(c) one can see that as the initial population of the $|\uparrow\rangle$ decreasing, the system is less likely to jump. The lower panels show the time-evolution of the components of the Bloch vector. Because the external driving is along $x$-direction, if the initial state lies in the $y$-$z$ plane the $\hat{\sigma}^{x}$ component is always zeros, as shown in Figs. 5(e) and (g).
@@
Figure 5: The dynamics of the driven single spin in Model I. The top panels show the time-evolution of the $K_{0}^{\o }(t)$ and $\sum_{\alpha}{K_{1}^{\alpha}}(t)$ with initial states $|\uparrow\rangle$ (a), $\frac{1}{\sqrt{2}}(|\uparrow\rangle+|\downarrow\rangle)$ (b) and $|\downarrow\rangle$ (c). The bottom panels show the time-evolution of the components of Bloch vectors. The initial states in (e) - (f) correspond to those in (a) - (c). The driving Rabi frequency is chosen as $\Omega=0.5$ and the decay rate is as given in Fig. 3.
@@ IV.2 Model II: Sudden death and revival of entanglement
V Comparison to the standard NMQJ method
In the approaches of jumplike unraveling the time-local master equation, the observable is computed through the ensemble average over all the possible states of the system. Because the probabilities of the states are interdependent in the non-Markovian regime, as shown in Eq. (40), computing these probabilities would be the most time-consuming part in the entire algorithm.
The standard NMQJ method is implemented by sampling a number of realizations with the identical initial (pure) state. In each realization, the system can undergo either the non-unitary evolution or normal/reversed quantum jumps. In order to approximate the true probabilities of the possible state, the number of realizations $N_{r}$ should be as large as possible. In the limit of $N_{r}\rightarrow\infty$, the existence probability in Eq. (13) is obtained. The improvement of our method lies in directly providing the existence probability. Moreover, one can choose an appropriate number $n^{*}$ to truncate the computation of the existence probabilities if the sum $\sum_{\alpha}{K_{n*}^{\alpha}}$ is below a threshold. For example, in Sec. IV, we have truncated at $n^{*}=2$ for both models. In order to show the validity of the truncation in the modified method and its advantage in reducing the computational time, we compare $\sum_{\alpha}{K_{2}^{\alpha}}$, which is the total existence probability of the trajectory class $\{H_{2}^{\alpha}\}$, obtained by our modified method and the standard NMQJ method with various $N_{r}$.
In Figs. 7(a)-(c), it is shown the time-evolution of the $\sum_{\alpha}{K_{2}^{\alpha}}$ for Model I, Model II in case (i) and case (ii), respectively. The continuous lines are computed according to the explicit expression while the discontinuous lines are computed by the standard NMQJ method with $N_{r}=10^{4}$, $10^{5}$ and $10^{6}$. As shown in Fig. 7(a), for the Model I the analytical result shows a good agreement with the results obtained by the standard NMQJ method. In particular, the value of $\sum_{\alpha}{K_{2}^{\alpha}}$ becomes zero at $t\approx 0.72$ as predicted by the analytical expression, although the peak of $N_{r}=10^{4}$ case is a little lower than the analytical result. As shown in Fig. 7(b), for the case (i) of Model II, the probability obtained with $N_{r}=10^{4}$ realizations deviates from the analytical result. As the $N_{r}$ increasing the probability converges to the analytical result gradually. Similar behavior can be observed for the case (ii) of Model II in Fig. 7(c).
Now we investigate the time consumption of the modified and standard NMQJ methods. In Fig. 7(d) the consumed time $t_{c}$ in the standard NMQJ method with various $N_{r}$ for both models are shown. The numerical computation is implemented on a regular computer workstation (Intel Xeon CPU E5-2680v3 and 64GB memory). As expected the time consumed in the computation grows dramatically as the $N_{r}$ increasing for the standard method. In order to obtain a more accurate result the $N_{r}$ should be larger than $10^{5}$ because there is still a non-zero tail at the end of the time-window for $N_{r}=10^{4}$. In this situation the $t_{c}$ is on the magnitudes of $10^{3}$ and $10^{5}$ seconds for $N_{r}=10^{5}$ and $10^{6}$, respectively, while the $t_{c}$ for the analytical result is on the magnitude $10\sim 100$.
In order to estimate the accuracy of the modified NMQJ method, we compute the Bures metric of the final state computed by the modified and standard NMQJ methods. In the modified NMQJ method, the final density matrix $\tilde{\rho}$ is reconstructed according to Eq. (14) with $n<2$. While in the standard NMQJ method, the final density matrix $\rho_{m}$ is obtained by averaging the states over all the $N_{r}$ realizations with $m=\log_{10}{N_{r}}$. The Bures metric of the state $\tilde{\rho}$ and $\rho_{m}$ is defined as follows,
$D_{B}(\tilde{\rho},\rho_{m})=\sqrt{2[1-F(\tilde{\rho},\rho_{m})]},$ (45)
with $F(\tilde{\rho},\rho_{m})=\text{tr}\sqrt{\sqrt{\tilde{\rho}}\rho_{m}\sqrt{\tilde{\rho}}}$ being the fidelity of two quantum states. The range of the Bures metric is $0\leq D_{B}\leq\sqrt{2}$. More specifically $D_{B}=0$ means two states are identical while $D_{b}=\sqrt{2}$ means two states are orthogonal. The Bures metric of $\tilde{\rho}$ and $\rho_{m}$ in Model I and both cases (i) and (ii) of Model II are listed in Tab. 1. One can see that the final state computed by the modified NMQJ method (with truncation $n^{*}=2$) shows good agreement with the one computed by the standard method. In particular, as expected, the Bures Metric approaches zero gradually as the $N_{r}$ increases.
We finish this section with a brief discussion on the storage resources. In the implementation of our modified NMQJ method, the state of each trajectory is stored in a $D_{H}\times 1$ column vector where $D_{H}$ is the dimension of the Hilbert space ($D_{H}=2$ for Model and $D_{H}=4$ for Model II). At the end of the period of positive decay rate, $N_{t}$ state vectors are stored for recording the information of each trajectory. The $N_{t}$ is number of discretized time points in the period of $\Delta(t)>0$. For examples, the $N_{t}=601$ in the models of this paper. Thus the storage space can be evaluated as $D_{H}\times N_{t}$. In contrast, the storage space needed in the standard NMQJ method is evaluated as $D_{H}\times N_{r}$. Therefore, with almost the same accuracy, the modified method costs less storage resource.
Figure 7: The total existence probability of the trajectory class $\{H_{2}^{\alpha}\}$ obtained by the modified NMQJ method and the standard NMQJ method with various realization number $N_{r}$ for (a) the Model I, (b) the case (i) of Model II, and (c) the case (ii) of Model II. The magenta solid line represents the result obtained from the modified method. The red dashed line, green dotted line, and blue dotted-dashed line represent for the results obtained from the standard method with $N_{r}=10^{4}$, $10^{5}$, and $10^{6}$, respectively. The consumed time $t_{c}$ (in the unit of seconds) of the simulation for each model are shown in panel (d). The blue square, red circle, and green triangle represents for the Model I, the case (i) and case (ii) of Model II, respectively. The horizontal axis of panel (d) denotes the realization number $N_{r}$. The consumed time of the modified method is plotted on the vertical axis of panel (d), indicated by the magenta arrow.
Table 1: The Bures metric of the final state computed by the modified NMQJ method and by the standard NMQJ method with various realization numbers $N_{r}$. The parameters of each case are chosen the same as in Figs. 7(a)-(c).
$N_{r}$ Model I Model II
($10^{m}$) case (i) case (ii)
$10^{4}$ 0.0152 0.0491 0.0499
$10^{5}$ 0.0138 0.0104 0.0345
$10^{6}$ 0.0162 0.0031 0.0130
VI Summary